# Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Text–Image Co-Editing

### v2

@@ A.2. Filters
Warm Warm tones (gold, amber, red, orange, yellow), high exposure, strong contrast $\rightarrow$ evoke happiness, comfort, nostalgia Emphasizes positivity, vitality, intimacy
Calm Cool tones (blue, green, purple) with balanced or lower saturation $\rightarrow$ convey calmness, wisdom, introspection Reflects contemplative and stable moods
@@
Dreamy Soft tones, lowered contrast, diffuse focus $\rightarrow$ suggest melancholy, intimacy, ethereality Supports subtle, nostalgic, introspective narration
Monochrome Removal of color, emphasis on light, shadow, texture $\rightarrow$ evoke nostalgia, timelessness, artistry Adopts reflective and universal tone
Table 4. Filter types with corresponding image and text effects.
A.3. LLM Prompts
Communication between different agents is carried out through JSON-formatted messages, ensuring efficient and reliable information transfer.
All prompts are structured in the following format:
  • Blue text indicates placeholders that will be replaced with task-specific information in different generation requests.
  • Pink text represents the mandatory output type of information generated by the agent.
The below is an example prompt for precise description generation of GPT-4o.
You are a visual story developer who analyzes screenshots to decode user visualization intent and creates detailed story segments that bring their creative vision to life. Process: 1) Examine the screenshot to understand what specific story content the user wants generated by identifying: printed text({text} ) which is the primary indicator of what the user wants you to generate, handwritten text expressing the user’s desired content direction and story focus where generation should address gaps and missing details, any images or illustrations with reference text {previous_text} for additional context, and hand-drawn sketches representing scenes from the user’s imagination. Synthesize these elements to understand the user’s envisioned story. The generated story should mainly focus on filling in content not covered in ({text} ) instead of still remain unknown. 2) If hand-drawn illustrations exist in the screenshot, return the information about the story scene conveyed in the illustration’s layout; if none exist, output ’none’. 3) Generate Focused Story Content: Using the existing written passages {full_text} only as background context to ensure logical consistency, create a NEW detailed story segment that elaborates on a specific scene or moment the user wants to visualize, focuses primarily on the intent expressed in the screenshot rather than expanding the existing text, contains concrete details, character emotions, environmental descriptions, dialogue, and interactions, maintains consistency with the global theme {global_theme}, and can reference previous text {previous_text} if relevant to the visualization goal. 4) The narrative should contain substantial plot or setting content, not just descriptive language. The generated story should introducing new, insightful elements based on the context and provide new direction of the story development that can reificate the story. The generated stories need to be imaginative with concrete content, not filled with uncertainties. Respond in JSON format: { "story": "A detailed paragraph of no larger than 100 words that creates NEW story content focused on the user’s screenshot intent, elaborating a specific scene with concrete details while maintaining logical consistency with the background context.", "intention": "The visualization intention read from the screenshot for story generation direction", "sketch_information":"Regarding line sketches, integrate them with story descriptions to capture the user’s envisioned layout and scene details communicated through hand-drawn imagery, directing the image generator to produce story scenes based on this layout guidance. Avoid generating stories that may trigger content moderation." }. Provide only the JSON response without markdown formatting or additional commentary. Let’s think step by step.

### v3

@@ A.4. Filters
Table 3. Filter types with corresponding image and text effects.
@@
Warm Warm tones (gold, amber, red, orange, yellow), high exposure, strong contrast $\rightarrow$ evoke happiness, comfort, nostalgia Emphasize positivity, vitality, intimacy
Calm Cool tones (blue, green, purple) with balanced or lower saturation $\rightarrow$ convey calmness, wisdom, introspection Reflect contemplative and stable moods
@@
Dreamy Soft tones, lowered contrast, diffuse focus $\rightarrow$ suggest melancholy, intimacy, ethereality Support subtle, nostalgic, introspective narration
Monochrome Removal of color, emphasis on light, shadow, texture $\rightarrow$ evoke nostalgia, timelessness, artistry Adopt reflective and universal tone
A.5. LLM Prompts
A.6. Codebook
Theme 1: Instrumental Interaction
   $\vdash$ instrumental operations
      $\vdash$ granularity control (using Lasso for detail extraction)
      $\vdash$ multimodal recombination (using Collage)
      $\vdash$ affective alignment (using Filters for tone)
      $\vdash$ perspective shift (viewpoint transformation)
      $\vdash$ open new narrative direction
Theme 2: Cognitive Process
   $\vdash$ externalization & traceability
      $\vdash$ visual history / story evolution
      $\vdash$ spatial organization (grouping)
      $\vdash$ visual checkpoints
   $\vdash$ cognitive offloading
      $\vdash$ offloading working memory
   $\vdash$ higher mental demand
      $\vdash$ not familiar with operations
      $\vdash$ substantial learning effort
Theme 3: Creative Support through Multimodality
   $\vdash$ bottom-up workflow
      $\vdash$ open-ended exploration
      $\vdash$ branching storylines
      $\vdash$ comparing modalities
      $\vdash$ divergent exploration
   $\vdash$ inspiration
      $\vdash$ serendipitous discovery (randomness as value)
      $\vdash$ perspective transformation
      $\vdash$ memory triggers
   $\vdash$ visual–text interaction
      $\vdash$ more vivid detailed description
      $\vdash$ sense of immersion
Theme 4: Ownership and Agency
   $\vdash$ control & ownership
      $\vdash$ resisting AI takeover
      $\vdash$ active curation
      $\vdash$ personal style alignment
   $\vdash$ metaphors of use
      $\vdash$ companion / sketchbook metaphor
      $\vdash$ co-pilot
      $\vdash$ free-exploration
Figure 14. The final coding tree. Main themes are marked in bold; sub-codes represent specific strategies and behaviors observed in the study.
The codebook tree of qualitative analysis